OpenStudy (anonymous):

Integrate sin^-1(x) dx please explain how to do it aswell

OpenStudy (anonymous):

there's a rule.

OpenStudy (anonymous):

integration by parts right?

OpenStudy (anonymous):

\[\int 1 \times sin^{-1}xdx\]

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

i dont understand it that well

OpenStudy (anonymous):

where does dv mean?

OpenStudy (anonymous):

v and u be functions of x \[\int uvdx = u\int vdx - \int [\int v dx \times \frac{d}{dx} u]\]

OpenStudy (anonymous):

so in this case u=1 and v=sin^-1(x)

OpenStudy (anonymous):

take u as \(sin^{-1}x\)

OpenStudy (anonymous):

sin ^ -1 = csc (x) the integral of that is.. - ln|csc x + cot x| + C

OpenStudy (anonymous):

does it make a difference which one is chosen as u?

OpenStudy (anonymous):

yes it does try it your self you will see

OpenStudy (anonymous):

integration by parts doesn't have x's in it idk what u guys are talking about haha

OpenStudy (anonymous):

ah I think he is talking about the inverse functions

OpenStudy (anonymous):

\[\int{sin^{-1}(x)dx = x*sin^{-1}+\sqrt{1 - x^2}}\]

OpenStudy (anonymous):

yeah thats the answer

OpenStudy (anonymous):

hmm Yes George is right

OpenStudy (anonymous):

this is really confusing lol

OpenStudy (anonymous):

I'm as confused as you are

OpenStudy (anonymous):

okay I will show you

OpenStudy (anonymous):

lol:)

OpenStudy (anonymous):

but yeah you use integration by parts to get the answer I wrote

OpenStudy (anonymous):

yes if you are talking about arcsin then mimi is right. if you are saying to the negative one power, look at mine.

OpenStudy (anonymous):

\[\int 1\times sin^{-1}x dx = sin^{-1}x \int 1 dx - \int [ \int 1 dx \times \frac{d (sin^{-1}x)}{dx}]\]

OpenStudy (aravindg):

ishaan help me

OpenStudy (aravindg):

ishaan

OpenStudy (anonymous):

ok thanks heaps guys:)

OpenStudy (anonymous):

\[= sin^{-1}x \times x - \int x \times \frac{1}{\sqrt{1-x^2}}\]

OpenStudy (anonymous):

Now integrate the second part

OpenStudy (anonymous):

\[t = 1 -x^2 \] \[dx = -\frac{dt}{x}\]

OpenStudy (anonymous):

\[dx = -\frac{dt}{2x}\]

OpenStudy (anonymous):

That gives you for the second part \[... + \frac{1}{2}\int \frac{1}{t^{\frac{1}{2}}}dt\]

OpenStudy (anonymous):

ah Now you can do

OpenStudy (anonymous):

ok thanks ill keep practicing with other questions thanks a lot for the help;)